# Spin-photon entanglement of a single Er3+ ion in the telecom band

### v1

@@
The rare earth ion (REI) Er3+ provides a direct spin-photon interface in the telecom band, but exploiting individual REIs as single photon sources and quantum memories is challenging because of their low photon emission rate and correspondingly increased sensitivity to decoherence. Recent work has demonstrated incorporating REIs into nanophotonic cavities to enhance the emission rate up to 1000-fold [30, 31, 32, 33, 34, 35, 36, 37], enabling single-shot readout [38, 39] and manipulating nearby nuclear spin ancillae [40, 41, 42], and remote entanglement of two Yb3+:YVO4 defects [14]. Careful selection and engineering of host materials has also led to improved coherence properties. In particular, Er3+:CaWO4 has demonstrated coherence times of 23 ms in small ensembles [43], and low enough spectral diffusion to enable indistinguishable single photon generation in a Si nanophotonic cavity [44]. However, despite these improvements, maintaining the joint spin and optical coherence during photon emission remains a challenge.

### v2

@@
The rare earth ion (REI) Er3+ provides a direct spin-photon interface in the telecom band, but exploiting individual REIs as single photon sources and quantum memories is challenging because of their low photon emission rate and correspondingly increased sensitivity to decoherence. Recent work has demonstrated incorporating REIs into nanophotonic cavities to enhance the emission rate up to 1000-fold [30, 31, 32, 33, 34, 35, 36, 37], enabling single-shot readout [38, 39, 40] and manipulating nearby nuclear spin ancillae [41, 42, 43], and remote entanglement of two Yb3+:YVO4 defects [14]. Careful selection and engineering of host materials has also led to improved coherence properties. In particular, Er3+:CaWO4 has demonstrated coherence times of 23 ms in small ensembles [44], and low enough spectral diffusion to enable indistinguishable single photon generation in a Si nanophotonic cavity [45]. However, despite these improvements, maintaining the joint spin and optical coherence during photon emission remains a challenge.